OpenStudy (anonymous):

How many terms are in the expansion (x + y)10?

OpenStudy (anonymous):

there is x and y so I guess two

OpenStudy (anonymous):

regardless if you foil out 10 you still have an x and y

OpenStudy (anonymous):

is it (x+y)^10

OpenStudy (anonymous):

No I'm pretty sure it's 12 because you add the 2 terms to 10

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

(x+y)^2 = x^2+2xy+y^2 <--- there are 3 terms here (x+y)^3 = x^3 + 3x^2y+3xy^2 + y^3 <--- there are 4 terms here (x+y)^4 = x^4+4x^3y + 6x^2y^2 + 4xy^3 + y^4 <--- there are 5 terms here etc etc See the pattern?